OpenStudy (ellie202000):

The length of a rectangle is 5 inches more than its width, x. The area of a rectangle can be represented by the equation x^2 + 5x = 300. What are the measures of the width and the length? Width= Length=

OpenStudy (anonymous):

So the area of the rectangle is 300

OpenStudy (anonymous):

which means the length is 300/x

OpenStudy (anonymous):

and the length is also x+5

OpenStudy (anonymous):

so 300/x=x+5 300=x2+5x

OpenStudy (anonymous):

which means x is 15

OpenStudy (ellie202000):

so would the equation look something like\[x ^{2}+5x-300=0\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I already solved that and x=15

OpenStudy (anonymous):

which means the width=15 and the length=20

OpenStudy (anonymous):



OpenStudy (anonymous):

when you realize you failed at making a picture

OpenStudy (anonymous):



OpenStudy (ellie202000):

Thank you so much!! That makes so much more sense. Do you think you would be able to help me with another problem?

OpenStudy (anonymous):

sure why not

OpenStudy (ellie202000):

The height, h, in feet of a golf ball above the ground after being hit into the air is given by the equation, \[h=-16t ^{2}+64t\]where t is the number of seconds elapsed since the ball was hit. How many seconds does it take for the golf ball to hit the ground?

OpenStudy (anonymous):

solve for the equation when h=0

OpenStudy (anonymous):

0=-16t2+64t t=4

OpenStudy (ellie202000):

Thank you so much again.